OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the given curves about the specified axis: y=x^2, y=0, x=2 about the y axis

OpenStudy (anonymous):

use shell method: V = 2pi ∫(radius) (height) V = 2pi ∫(x) (x²) dx V = 2pi ∫x^3 dx V = (pi/2)x^4 (evaluate from 0 to 2) V = pi/8

OpenStudy (anonymous):

I meant *8pi*